radar detected an unidentified plane 5000 km away, approaching at 700 kph. Fifteen minutes later, an interceptor plane was dispatched, travelling at 800 kph. How long did the interceptor take tp reach the approaching plane?
Answer by mananth(16946)


radar detected an unidentified plane 5000 km away, approaching at 700 kph.
In 1/4 hour the plane wouldbe nearer by d= speed * time
700*1/4 = 175 km
5000-175 = 4825
Fifteen minutes later, = 1/4 hour
an interceptor plane was dispatched, travelling at 800 kph.
Both planes approaching each other.
Effective speed = 700+800 =1500 kph
Distance between them = 4825 km
time = distance / speed
= 4825/1500
= 3.22 hours



the interceptor take 3.22 hours to reach the approaching plane